OpenStudy (anonymous):

Given y=x^(2)-8x+9, Complete the square and convert the equation to y=a(x-h)^(2)+k

OpenStudy (helder_edwin):

you know this \[ \large (a+b)^2=a^2+2ab+b^2 \]

OpenStudy (helder_edwin):

know try finding similarities with what you have a^2 corresponds with x^2 right?

OpenStudy (anonymous):

yea.

OpenStudy (helder_edwin):

now look at the middle term. a corresponds to x then 2b=-8 agree?

OpenStudy (anonymous):

yes.

OpenStudy (helder_edwin):

what is the value of b?

OpenStudy (anonymous):

-4?

OpenStudy (helder_edwin):

yes then b^2=

OpenStudy (anonymous):

16

OpenStudy (helder_edwin):

great do you have 16?

OpenStudy (helder_edwin):

are you there?

OpenStudy (anonymous):

as one of the answers. no. The answers are A. y=(x-4)^2+7 B.y=(x-4)^2-7 C.y=(x-8)^2+9 D.y=(x-4)^2+9 E.None of the above

OpenStudy (helder_edwin):

wait now you have \[ \large y=x^2-8x+9 \] you add and substract what you found b=16 \[ \large y=x^2-8x+16+9-16 \] \[ \large y=(x-4)^2+ \]

OpenStudy (helder_edwin):

you tell me which one is the correct

OpenStudy (anonymous):

9

OpenStudy (helder_edwin):

9-16=

OpenStudy (anonymous):

-7 so the answer is B?

OpenStudy (helder_edwin):

yes! great work did u understand?

OpenStudy (anonymous):

yes i did. thank you so much.!

OpenStudy (helder_edwin):

u r welcome